OpenStudy (anonymous):

Solve : 3^(4x-7) = 4^(2x+3)

OpenStudy (rulnick):

(4x-7) log 3 = (2x+3) log 4 (4log3-2log4)x=3log4+7log3 x=(3log4+7log3)/(4log3-2log4) x=(6log2+7log3)/(4(log3-log2))

OpenStudy (rulnick):

... or about 7.306.